OpenStudy (anonymous):

a 3m x 3m x 4m room contains air with a moisture content of 11.6 gram H20/dry air at 1 atm (absolute) pressure and 25 degree Celsius. the dry air component can be assumed to have a composition of 21 mol% oxygen and 79 mol% nitrogen. 1) calculate the mass fraction of water vapour in the room.2 2) mole fraction of water vapour in the room?

OpenStudy (cuanchi):

moisture content of 11.6 g h20/dry air =moisture content of 11.6 g H2O % dry air ???

OpenStudy (anonymous):

any one there to solve above question?

OpenStudy (cuanchi):

please tell me what do you mean with this, I can not help you without that! "11.6 g h20/dry air"

OpenStudy (anonymous):

it means 11.6 gram H2O/kg dry air

OpenStudy (cuanchi):

molar mass of dry air (0.21*32+0.79*28)= 28.84g/mol 11.6 g H2O /18g = 0.64 mol H2O/kg dry air 1000g dry air/28.84g =34.67 mol dry air 11.6 g H2O/kg of dry air = 0.64 mol H2O/34.67 mole Dry air= 1mol H2O/54 mol dry air = 54mol dry air/mol of H2O 0.0185 mol H2O/mol of dry air you have 1mol of H2O every 54 mole of dry air or every 55 mole of gas total volume of your room the total volume of your room is 3*3*4= 36 cubic meters= 36000L the molar volume of a ideal gas in STP conditions is 22.4L then you can calculate how many moles of gas you have in the room 36000L/22.4L= 1607.14 moles of gas for every 55 moles of gas one is H2O 1607.14/55= 29.22 moles of H2O in the room => 29.22*18= 252.97 g H2O 29.2 moles of H2O in the room = 29.22*18 = 252.97 g H2O 1576.8 moles of dry air in the room = 1576.8*28.84= 45474.9 g dry air 1607 moles of gas in the room =45727.882 g 45727.882/1607 =28.46g/mol gas in the room